My Notes Ask Your Teacher Values of modulus of elasticity (MOE, the ratio of stress, .e., force per unit area, to strain, .e.I NEED A SOLUTION ASAP PLEASE
Answer #1

b)

X Value=   40      
Confidence Level=   95%      
          
          
Sample Size , n=   27      
Degrees of Freedom,df=n-2 =   25      
critical t Value=tα/2 =   2.060   [excel function: =t.inv.2t(α/2,df) ]  

standard error, S(ŷ)=    0.179              
margin of error,E=t*Std error=t* S(ŷ) =   2.0595   *   0.1789   =   0.3684
                  
Confidence Lower Limit=Ŷ +E =    7.555   -   0.3684   =   7.187
Confidence Upper Limit=Ŷ +E =   7.555   +   0.3684   =   7.923

c)

X̅ =    45.11
Σ(x-x̅)² =Sxx   4582.9
Standard Error of the Estimate,Se=   0.87

standard error, S(ŷ)=Se*√(1+1/n+(X-X̅)²/Sxx) =   0.8835              
margin of error,E=t*std error=t*S(ŷ)=    2.0595   *   0.88   =   1.8196
                  
Prediction Interval Lower Limit=Ŷ -E =   7.555   -   1.8196   =   5.735
Prediction Interval Upper Limit=Ŷ +E =   7.555   +   1.8196   =   9.375

d)

the simultaneous confidence interval is at least 100(1-2α) = 100(1-2*0.05) = 90%